OpenStudy (anonymous):

Write the equation of the line that contains the indicated pair of points. Express the final equation in standard form. (-5, 4), (3, -3)

OpenStudy (tamtoan):

with that 2 points, what is the slope (m) of the line?

OpenStudy (anonymous):

I got -7/8 ?

OpenStudy (tamtoan):

ok the standard form of the equation is y = mx + b ...now you only need to find the value of b ...and since the line will pass by that 2 points..so use 1 point to find b...what do you get for b ?

OpenStudy (anonymous):

one sec..

OpenStudy (anonymous):

I'm stuck here .... y - 4 = -7/8(x + 5)

OpenStudy (tamtoan):

your equation is y = mx + b ...since you already have m = -7/8, so your equation now is y = -7/8x + b ...now pick a point and substitute value of (x,y) of that point into the equation, then find b

OpenStudy (anonymous):

Ok.. If I subsititue 8 for x.. I should get y = -7 + b

OpenStudy (anonymous):

..I'm confused

OpenStudy (tamtoan):

no, of the 2 points you have (-5,4) and (3,-3) ...each value of x there yield 1 value of y...like when x = -5, y = 4 ...sub that into the equation and find b

OpenStudy (tamtoan):

did you get value of b yet?